OpenStudy (anonymous):

Write an equation of the line that is parallel to -x + y = 5 and passes through the point (2, -5).

OpenStudy (campbell_st):

the slope of the parallel line is m = -1 the slope intercept form of the line is y = mx + b you have x = 2, y = -5 and m = -1 substitute and find b, then you'll get the equation of the line

OpenStudy (anonymous):

oooo oh ok thank you campbell

OpenStudy (campbell_st):

oops... the slope is m = 1 the slope intercept form of the given line is y = x + 5 so m = 1 sorry for the typo...

OpenStudy (anonymous):

alright and why wouldnt it be x - 3

OpenStudy (campbell_st):

well -5 = 1*2 + b b = -7 so y = x - 7 that should be the answer

OpenStudy (anonymous):

ok thanks